# FedDyMem: Efficient Federated Learning with Dynamic Memory and Memory-Reduce for Unsupervised Image Anomaly Detection

### v1

@@ IV-C Computation of Local Metric Loss: $\ell({{\cal M}^{n,i,t}},{{\cal M^{\prime}}^{n,t-1}})$
7:     ${\{{\cal M}^{n,i,0}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}\leftarrow[{\cal M}^{0},{\cal M}^{1},\cdots]$;
8:     ${\cal M}^{n,0}_{reduce}\leftarrow$ Compute by Algorithm 2 with $n$, $t=0$ and ${\{{\cal M}^{n,i,0}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}$;
9:     ${\bf W}^{{\cal P},n,0}\leftarrow{\bf W}^{\cal P}$, ${\bf W}^{{\cal G},n,0}\leftarrow{\bf W}^{\cal G}$, ${{\bf G}^{n,0}}\leftarrow{\bf G}$
10:  end for
11:  ${\{{\cal M}^{n,0}_{reduce}\}_{n=0}^{N}}\leftarrow[{\cal M}^{0,0}_{reduce},{\cal M}^{1,0}_{reduce},\cdots]$;
12:  ¯0superscript¯$\bar{\cal M}^{0}\leftarrow$ Compute by Algorithm 3 with $t=0$ and ${\{{\cal M}^{n,0}_{reduce}\}_{n=0}^{N}}$;
13:  for $n=0$ to $N$ do
14:     ${\cal M^{\prime}}^{n,0}\leftarrow\bar{\cal M}^{0}$;
15:  end for
16:  return  ${\{{\cal M^{\prime}}^{n,0}\}}_{n=0}^{N}\leftarrow[{\cal M^{\prime}}^{0,0},{\cal M^{\prime}}^{1,0},\cdots]$
IV-E ¯$\bar{\cal M}^{t}\stackrel{{\scriptstyle agg}}{{\longleftarrow}}{\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}$ on Server
At round $t$, the server $\cal S$ collects the newly generated memory banks from $N$ clients, denoted as ${\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}\in{{\mathbb{R}^{N\times H\times W\times C}}}$. The objective of the server is to aggregate these memory banks into a unified, representative memory bank ¯$\bar{\cal M}^{t}$. This process is formalized as ${\cal S}(\cdot,\cdot)\colon{{\mathbb{R}}^{N\times H\times W\times C}}\to{{\mathbb{R}}^{H\times W\times C}}$. Representation-based federated learning approaches, such as those proposed in [34, 63], utilize weighted aggregation of class prototypes during model updates. However, significant feature bias in the early training stages leads to aggregated memory that deviates significantly from the client-specific memory, which often contains incomplete product types. This misalignment hinders convergence, posing a challenge for clients with limited data diversity. Thus, we employ a clustering-based approach to aggregate the collected memory banks effectively. Specifically, the memory features are represented as $N\times H\times W$ patch features, denoted by ${m_{(h,w)}^{n}}\in{\mathbb{R}^{C}}$. These patch features are subsequently clustered using the $K$-means algorithm, where the number of clusters, $K$, corresponds to $H\times W$. The clustering results indicate the overall distribution and similarity relationships between the patch-level features stored across all banks at the current stage of the process. Cluster centers describe the representative features of a memory bank. Let $c_{k}\in{\mathbb{R}^{C}}$ denote the center of the $k$-th cluster, serving as a global memory bank updated by ¯$\bar{\cal M}^{t}={f_{resize}}(\{c_{k}\}_{k=0}^{K},(H,W,C))$ and shared across individual clients, $\forall n\in N,\ {\cal M^{\prime}}^{n,t}=\bar{\cal M}^{t}$. The aggregation methodology is outlined in Algorithm 3, while the comprehensive training procedure for FedDyMem is presented in Algorithm 5.
Algorithm 5 Training for FedDyMem.
0:  $N$ Clients $\{{\cal C}_{n}\}_{n=1}^{N}$ with ${\{{\cal D}_{n}^{train}\}}_{n=0}^{N}$, one server $\cal S$, the number of rounds $T$, the local epochs $E$.
1:  ${\{{\cal M^{\prime}}^{n,0}\}}_{n=0}^{N}\leftarrow$ Initialization by Algorithm 4;
2:  for $t=1$ to $T$ do
3:     for each client $n$ in parallel do
4:        ${\cal M}^{n,t}_{reduce}\leftarrow$ ClientUpdate(n, t, E, ${\cal D}_{n}^{train}$);
5:     end for
6:     ${\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}\leftarrow$ Collection all memory bank from $N$ clients;
7:     ¯$\bar{\cal M}^{t}\leftarrow$ Upload ${\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}}$ to $\cal S$ and computed by Algorithm 3;
8:  end for
9:  Save all well trained {{\${\bf W}^{{\cal P},n,T}$, ${\bf W}^{{\cal G},n,T}$, ${{\bf G}^{n,T}}\}_{n=0}^{N}$ and ¯$\bar{\cal M}^{T}$ locally.
9:
10:  for $e=0$ to $E$ do
11:     for $i=0$ to $\left|{\cal D}_{n}^{train}\right|$ do
12:        ${\cal M}^{i}\leftarrow$ Compute by Algorithm 1 with $t$, $n$, $x_{i}^{n}$;
13:        $\ell\leftarrow$ Compute loss by Equation (12) with ${\cal M}^{i}$ and ${\cal M^{\prime}}^{n,t-1}$;
14:        Update {{\${\bf W}^{{\cal P},n,t}$, ${\bf W}^{{\cal G},n,t}$, ${{\bf G}^{n,t}}\}$ by $\ell$ and Adam Optimizer [64];
15:     end for
16:  end for
17:  ${\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}\leftarrow[{\cal M}^{0},{\cal M}^{1},\cdots]$;
18:  ${\cal M}^{n,t}_{reduce}\leftarrow$ Compute by Algorithm 2 with $n$, $t$ and ${\{{\cal M}^{n,i,t}\}_{i=0}^{\left|{\cal D}_{n}^{train}\right|}}$;
19:  return  ${\cal M}^{n,t}_{reduce}$
IV-F Privacy Preserving Discussion
FedDyMem facilitates the exchange of statistical information, i.e., memory banks, between the clients and the server, instead of transmitting model parameters, thereby enhancing privacy protection [34]. The memory features that are transmitted are derived from multi-layer low-dimension representations, which is an irreversible process[65]. In practice, additional privacy-preserving techniques can be incorporated into the FedDyMem framework to further improve system reliability [66] and ensure robust data protection [67].
V Convergence Analysis
We provide the convergence analysis for FedDyMem considering the non-convex nature of its loss function. Some common assumptions based on existing frameworks are reasonably presented below:
Assumption 1 (Piecewise $L_{2}$-Lipschitz Smooth). Let ${\rm{dis}}(m_{(h,w)},m_{k}^{\prime})={th}$ denote the discontinuity point. For any $t,t^{\prime}>0$, there exists $L>0$ satisfying:
$\|\nabla\ell(\mathcal{M}^{n,i,t},\mathcal{M^{\prime}}^{n,t-1})-\nabla\ell(\mathcal{M}^{n,i,t^{\prime}},\mathcal{M^{\prime}}^{n,t^{\prime}-1})\|_{2}$ (17)
$\leq L\|\mathcal{M}^{n,i,t}-\mathcal{M}^{n,i,t^{\prime}}\|_{2}.$
Assumption 2 (Bounded Gradients). The gradients of the loss function are bounded, i.e., there exists $G>0$ satisfying:
$\|\nabla\ell(\mathcal{M}^{n,i,t},\mathcal{M^{\prime}}^{n,t-1})\|_{2}\leq G.$ (18)
Assumption 3 (Unbiased Gradient and Bounded Variance). The stochastic gradient $g_{t}$ is an unbiased estimate of the true gradient, i.e., $\mathbb{E}[g_{t}]=\nabla\ell(\mathcal{M}^{n,i,t},\mathcal{M^{\prime}}^{n,t-1})$. Furthermore, the variance of $g_{t}$ is bounded, i.e., there exists $\sigma^{2}>0$ satisfying:
$\mathbb{E}[\|g_{t}-\nabla\ell(\mathcal{M}^{n,i,t},\mathcal{M^{\prime}}^{n,t-1})\|_{2}^{2}]\leq\sigma^{2}.$ (19)
Assumption 4. Based on the loss function’s structure, the hyperparameter $th$ is constrained as:
$th\in\left[\min_{h,w,k}{\rm{dis}}(m_{(h,w)},m_{k}^{\prime}),\frac{1}{HWK}\sum_{h,w,k}^{H,W,K}{\rm{dis}}(m_{(h,w)},m_{k}^{\prime})\right]$
, ensuring the loss function $\ell$ remains finite and satisfies the operational properties in Assumptions 1-4.
Theorem 1 (Loss Reduction). For any client, after every communication round, when the learning rate satisfies $0<\eta<\frac{2}{L}$, and letting $\mathcal{M}^{t}=\mathcal{M}^{n,i,t},\mathcal{M^{\prime}}=\mathcal{M^{\prime}}^{n,t-1}$, the expected loss after $T$ iterations satisfies:
$\mathbb{E}[\ell(\mathcal{M}^{T},\mathcal{M^{\prime}})]$ $\leq\ell(\mathcal{M}^{n,i,1},\mathcal{M^{\prime}}^{n,0})+\frac{T\eta^{2}L\sigma^{2}}{2}$ (20)
@@ V Convergence Analysis
$\quad-\left(\eta-\frac{\eta^{2}L}{2}\right)\sum_{t=1}^{T-1}\mathbb{E}[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}].$
Proof. For each iteration, based on the Assumptions above, the loss function satisfies:
$\mathbb{E}[\ell(\mathcal{M}^{t+1},\mathcal{M^{\prime}})]$ $\leq\mathbb{E}[\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})]-\eta\mathbb{E}[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}]$
$\quad+\frac{\eta^{2}L}{2}\mathbb{E}[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}]$
$\leq\mathbb{E}[\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})]+\frac{\eta^{2}L\sigma^{2}}{2}$
@@
$\quad-\left(\eta-\frac{\eta^{2}L}{2}\right)\mathbb{E}[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}].$
After $T$ iterations, we obtain:
$\mathbb{E}[\ell(\mathcal{M}^{T},\mathcal{M^{\prime}})]$ $\leq\ell(\mathcal{M}^{n,i,1},\mathcal{M^{\prime}}^{n,0})+\frac{T\eta^{2}L\sigma^{2}}{2}$
$\quad-\left(\eta-\frac{\eta^{2}L}{2}\right)\sum_{t=1}^{T-1}\mathbb{E}[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}].$
Theorem 1 expresses that by appropriately selecting $\eta$ and $\sigma$, the penalty term $\frac{T\eta^{2}L\sigma^{2}}{2}$ in the loss descent inequality can be controlled, such that the term $-\left(\eta-\frac{\eta^{2}L}{2}\right)\sum_{t=1}^{T-1}\mathbb{E}\left[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}\right]$ plays a dominant role in the gradient descent of the loss function. In the subsequent Theorem 2, we provide the convergence results of the model.
Theorem 2 (FedDyMem Convergence). Based on the previous assumptions and Theorem 1, when $th$ satisfies the constraint in Assumption 4, for any $\epsilon>\frac{\eta L\sigma^{2}}{2(\eta-\frac{\eta^{2}L}{2})}$, there exists:
$T=\frac{\ell(\mathcal{M}^{n,i,1},\mathcal{M^{\prime}}^{n,0})-\underline{\ell}}{\epsilon\left(\eta-\frac{\eta^{2}L}{2}\right)},$
such that for $t>T$,
$\frac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}[\|\nabla\ell(\mathcal{M}^{n,i,t},\mathcal{M^{\prime}}^{n,t-1})\|_{2}^{2}]\leq\epsilon.$ (21)
Proof. From Theorem 1, we have:
$\mathbb{E}[\ell(\mathcal{M}^{T},\mathcal{M^{\prime}})]$ $\leq\ell(\mathcal{M}^{n,i,1},\mathcal{M^{\prime}}^{n,0})+\frac{T\eta^{2}L\sigma^{2}}{2}$
$\quad-\left(\eta-\frac{\eta^{2}L}{2}\right)\sum_{t=1}^{T-1}\mathbb{E}[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}].$
Since $\ell(\mathcal{M},\mathcal{M^{\prime}})$ has a non-negative lower bound $\underline{\ell}$, satisfying $0<\underline{\ell}<\ell$, we write:
$\underline{\ell}$ $\leq\ell(\mathcal{M}^{T},\mathcal{M^{\prime}})$
$\leq\ell(\mathcal{M}^{n,i,1},\mathcal{M^{\prime}}^{n,0})+\frac{T\eta^{2}L\sigma^{2}}{2}$
$\quad-\left(\eta-\frac{\eta^{2}L}{2}\right)\sum_{t=1}^{T-1}\mathbb{E}[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}].$
After rearranging and simplifying:
$\frac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}[\|\nabla\ell(\mathcal{M}^{t},\mathcal{M^{\prime}})\|_{2}^{2}]\leq\frac{\ell(\mathcal{M}^{n,i,1},\mathcal{M^{\prime}}^{n,0})-\underline{\ell}}{T(\eta-\frac{\eta^{2}L}{2})}+\frac{\eta L\sigma^{2}}{2(\eta-\frac{\eta^{2}L}{2})}.$
Letting this upper bound be less than or equal to $\epsilon$, we solve for $T$:
$T\geq\frac{\ell(\mathcal{M}^{n,i,1},\mathcal{M^{\prime}}^{n,0})-\underline{\ell}}{\epsilon(\eta-\frac{\eta^{2}L}{2})}.$
VI Experiments

### v2

@@ 4.3 Computation of Local Metric Loss: $\ell({{\cal M}^{n,i,t}},{{\cal M^{\prime}}^{n,t-1}})$
7:  $\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}$$\leftarrow$ Collection all memory bank from NN clients;
8:  $\bar{\cal M}^{t}\leftarrow$ Upload $\{{\cal M}^{n,t}_{reduce}\}_{n=0}^{N}$} to $\cal S$ and computed by Algorithm 3;
9:end for
10:Save all well trained ${\bf W}^{{\cal P},n,T}$, ${\bf W}^{{\cal G},n,T}$, ${{\bf G}^{n,T}}\}_{n=0}^{N}$ and $\bar{\cal M}^{T}$ locally.
11:
12:for $e=0$ to $E$ do
13:  for $i=0$ to $\left|{\cal D}_{n}^{train}\right|$ do
14:   ${\cal M}^{i}\leftarrow$ Compute by Algorithm 1 with $t$, $n$, $x_{i}^{n}$;
15:   $\ell\leftarrow$ Compute loss by Equation (12) with ${\cal M}^{i}$ and ${\cal M^{\prime}}^{n,t-1}$;
16:   Update ${\bf W}^{{\cal P},n,t}$, ${\bf W}^{{\cal G},n,t}$, ${\bf G}^{n,t}$}\} by $\ell$ and Adam Optimizer;
17:  end for
18:end for
[…]
21:return ${\cal M}^{n,t}_{reduce}$
4.6 Testing for Anomaly Detection on Clients
During the testing phase, as shown in Fig. 2(b), each client employs the globally aggregated memory bank $\bar{\cal M}^{T}$, received from the server at the end of training, to perform anomaly detection on its local test samples. Following prior works ([roth2022towards, lee2022cfa]), we utilize a nearest neighbor search in the feature space to calculate the anomaly score for each test sample $x^{test}\in{{\cal D}^{test}}$.
Given a test image $x^{test}$ on $n$-th client, the well trained local models ${\bf W}^{{\cal P},n,T}$, ${\bf W}^{{\cal G},n,T}$ and ${\bf G}^{n,T}$ are used to extract the memory feature representation ${\cal M}^{test}\in\mathbb{R}^{H\times W\times C}$ following the same procedure as in training. For each spatial location $(h,w)$, the corresponding feature patch $m^{test}_{(h,w)}\in\mathbb{R}^{C}$ is compared with the final global memory bank $\bar{\cal M}^{T}=\{\bar{m}_{d}\in\mathbb{R}^{C}\}_{d=1}^{H\times W}$ to compute its anomaly score. The pixel-wise anomaly score $a_{(h,w)}$ is computed as the minimum Euclidean distance to its $K$ nearest neighbors in the memory bank. And the image-level anomaly score ${\cal A}^{test}$} is computed by the softmax-weighted maximum strategy:
${{\cal A}^{test}}=\max_{(h,w)}\left(a_{(h,w)}\cdot\frac{\exp(a_{(h,w)})}{\sum_{h^{\prime},w^{\prime}}\exp(a_{(h^{\prime},w^{\prime})})}\right).$ (17)
4.7 Privacy Analysis
In federated learning, Mutual Information(MI) is a fundamental tool for evaluating privacy. Prior studies ([zhang2023fedcr, alsulaimawi2024federated, ji2025re, tan2024defending]) indicate that MI provides a principled paradigm for privacy analysis and can jointly improve privacy and utility. Consistent with prior work, we adopt a MI–based methodology to analyze FedDyMem’s privacy. For two statistics $X$ and $Y$, their mutual information $I(X;Y)$ quantifies statistical dependence; under Gaussian assumptions,
$I(X;Y)=-\tfrac{1}{2}\log\!\big(1-\rho_{xy}^{2}\big).$ (18)
where $\rho_{xy}$ denotes the Pearson correlation coefficient. A smaller $I(X;Y)$ implies that an adversary can extract less information about $X$ from observing $Y$, indicating stronger privacy.
In our FedDyMem, let $x^{j}$ be a raw sample, ${\cal M}^{j}=f(x^{j})$ denotes its corresponding memory feature. The reduced memory ${\cal M}_{reduce}\leftarrow[{\cal M}^{0},{\cal M}^{1},\cdots]$ is obtained via weighted averaging and the rEMA update, which induces the Markov chain $x^{j}\to({\cal M}^{0},{\cal M}^{1},\cdots)\to{\cal M}_{reduce}$. By the data–processing inequality (DPI), the mutual information decreases after memory-reduce, i.e.,
$I(x^{j};{\cal M}_{reduce})\leq I(x^{j};{\cal M}^{j}),$ (19)
which qualitatively captures the privacy effect of the memory–reduce module. Building on this Lemma 1 together with Theorem 1 provides a more rigorous and quantitative formulation of this privacy–preserving capability.
Lemma 1 (Correlation Quantitative Bound).
Let $\{Y_{i}\}_{i=1}^{N}$ be scalar statistics and define the normalized weighted average $Y=\sum_{i=1}^{N}\tilde{w}_{i}Y_{i}$ with $\tilde{w}_{i}\geq 0$, $\sum_{i}\tilde{w}_{i}=1$. Assume:
  1. 1.
    $X_{j}$ is independent of $Y_{i}$ for $i\neq j$.
  2. 2.
    Each $Y_{i}$ has finite, strictly positive variance $\sigma_{Y_{i}}^{2}$, and $\mathrm{Cov}(Y_{i},Y_{k})\geq 0$ for $i\neq k$.
Denote $\sigma_{Y_{i}}^{2}=\mathrm{Var}(Y_{i})$, then
$|\rho(X_{j},Y)|\;\leq\;\frac{\tilde{w}_{j}\,\sigma_{Y_{j}}}{\sqrt{\sum_{i=1}^{N}\tilde{w}_{i}^{2}\,\sigma_{Y_{i}}^{2}}}\,|\rho(X_{j},Y_{j})|.$ (20)
Proof.
By linearity and independence, $\mathrm{Cov}(X_{j},Y)=\tilde{w}_{j}\,\mathrm{Cov}(X_{j},Y_{j})$, hence
$|\rho(X_{j},Y)|=\frac{|\mathrm{Cov}(X_{j},Y)|}{\sigma_{X_{j}}\,\sigma_{Y}}=\frac{\tilde{w}_{j}\,|\mathrm{Cov}(X_{j},Y_{j})|}{\sigma_{X_{j}}\,\sigma_{Y}}.$ (21)
For the denominator,
$\mathrm{Var}(Y)=\sum_{i=1}^{N}\tilde{w}_{i}^{2}\,\sigma_{Y_{i}}^{2}+2\sum_{1\leq i<k\leq N}\tilde{w}_{i}\tilde{w}_{k}\,\mathrm{Cov}(Y_{i},Y_{k}).$ (22)
Since $\mathrm{Cov}(Y_{i},Y_{k})\geq 0$, with $\sigma_{Y}=\sqrt{\mathrm{Var}(Y)}$, the variance satisfies $\sigma_{Y}=\sqrt{\mathrm{Var}(Y)}\geq\sqrt{\sum_{i=1}^{N}\tilde{w}_{i}^{2}\,\sigma_{Y_{i}}^{2}}$.
Substituting this bound into the denominator of (21) with $|\mathrm{Cov}(X_{j},Y_{j})|=\sigma_{X_{j}}\sigma_{Y_{j}}|\rho(X_{j},Y_{j})|$, completes the proof. ∎
Theorem 1 (Memory-Reduce Privacy).
For any client sample $x^{j}$ with memory feature ${\cal M}^{j}$ and reduced memory ${\cal M}_{reduce}$,
$I(x^{j};{\cal M}_{reduce})\ll I(x^{j};{\cal M}^{j}).$ (23)
Proof.
Following (14)(16) we define $x^{j}$ as $X_{j}$ and ${\cal M}_{reduce}$ as the normalized weighted average $Y$ in Lemma 1. Thus, the correlation between $x^{j}$ and the reduced memory satisfies
$|\rho(x^{j},{\cal M}_{reduce})|\leq\frac{\tilde{w}_{j}\sigma_{{\cal M}^{j}}}{\sqrt{\sum_{i=1}^{|{\cal D}|}\tilde{w}_{i}^{2}\sigma_{{\cal M}^{j}}^{2}}}|\rho(x^{j},{\cal M}^{j})|,$ (24)
where $|{\cal D}|$ denotes the number of local samples at the client.
Since mutual information is monotone in $|\rho|$, we obtain
$\frac{I(x^{j};{\cal M}_{reduce})}{I(x^{j};{\cal M}^{j})}\leq\frac{-\tfrac{1}{2}\log\!\Big(1-\frac{\tilde{w}_{j}^{2}\sigma_{{\cal M}^{j}}^{2}}{\sum_{i=1}^{|{\cal D}|}\tilde{w}_{i}^{2}\sigma_{{\cal M}^{j}}^{2}}\rho(x^{j},{\cal M}^{j})^{2}\Big)}{-\tfrac{1}{2}\log\!\big(1-\rho(x^{j},{\cal M}^{j})^{2}\big)}.$ (25)
Moreover, since the averaging denominator grows with the number of samples $|{\cal D}|$, the effective scaling factor
$\frac{\tilde{w}_{j}^{2}\sigma_{{\cal M}^{j}}^{2}}{\sum_{i=1}^{|{\cal D}|}\tilde{w}_{i}^{2}\sigma_{{\cal M}^{j}}^{2}}\;\ll\;1,$ (26)
which proves the Theorem 1. ∎
By (26), the per-sample scaling vanishes as $|{\cal D}|$ grows. Under the bounded-variance conditions in Lemma 1, this yields $|\rho(x^{j},{\cal M}_{reduce})|\to 0$ and therefore $I(x^{j};{\cal M}_{reduce})$} as $|{\cal D}|\to\infty$, making the per-sample leakage asymptotically negligible. This establishes that the privacy leakage diminishes both with more samples, providing a strong quantitative guarantee, confirming that the proposed memory-reduce is essential for privacy preservation. Furthermore, FedDyMem exchanges statistical summaries (memory banks) rather than model parameters ([tan2022fedproto]) and the transmitted memory features are derived from multi-layer, low-dimensional representations via irreversible mappings ([li2023prototype]).
5 Convergence Analysis
We analyze the convergence of FedDyMem under a non-convex loss. Let the loss at local iteration $e$ in communication round $t$ be $\mathcal{L}_{t,e}:=\mathbb{E}_{x\sim\mathcal{D}_{n}}\!\big[\ell(\mathcal{M}^{t,e},\ \mathcal{M}^{\prime\,t-1})\big]$, where $\mathcal{M}^{t,e}=f(x;\theta_{t,e})$ and its Jacobian as $J_{f}(x;\theta_{t,e})$. Here $\mathcal{M}^{\prime\,t-1}$ is the center memory received at round $t$. Some common assumptions based on existing frameworks are reasonably presented below.
Assumption 1 ($L_{M}$-smooth).
Within a fixed communication round $t$, treat $\mathcal{M}^{\prime\,t-1}$ as a constant. Excluding the measure-zero non-differentiable set,
$\|\nabla_{\mathcal{M}}\ell(\mathcal{M},\mathcal{M}^{\prime})-\nabla_{\mathcal{M}}\ell(\tilde{\mathcal{M}},\mathcal{M}^{\prime})\|_{2}\leq L_{M}\|\mathcal{M}-\tilde{\mathcal{M}}\|_{2},$ (27)
and $\|\nabla_{\mathcal{M}}\ell(\mathcal{M},\mathcal{M}^{\prime})\|_{2}\leq G_{M}$. By the chain rule, the composite smoothness constant satisfies $L_{\theta}\leq B_{J}^{2}L_{M}$, and $\|\nabla_{\theta}\ell(\mathcal{M},\mathcal{M}^{\prime})\|_{2}\leq B_{J}G_{M}$.
Assumption 2 (Bounded features and Jacobian).
There exist constants $R_{M},B_{J}>0$ such that, for any $\theta,x$, satisfies $\|f(x;\theta)\|_{2}\leq R_{M}$, $\|J_{f}(x;\theta)\|_{2}\leq B_{J}$.
Assumption 3 (Unbiased stochastic gradients with bounded variance).
The local SGD gradient estimate $g_{t,e}$ satisfies
$\mathbb{E}\big[g_{t,e}\big]=\nabla_{\theta}\,\mathbb{E}_{x\sim\mathcal{D}_{n}}\!\big[\ell(\mathcal{M}^{t,e},\mathcal{M}^{\prime\,t-1})\big],$ (28)
and $\mathbb{E}\big\|g_{t,e}-\nabla_{\theta}\mathbb{E}_{x}[\ell(\mathcal{M}^{t,e},\mathcal{M}^{\prime\,t-1})]\big\|_{2}^{2}\ \leq\ \sigma^{2}.$
Assumption 4 (Lipschitz Continuity in $\mathcal{M}^{\prime}$).
Client-side reduction and server aggregation are convex combinations. Together with Assumption 2, this implies $\|\mathcal{M}^{\prime\,t}\|_{2}\leq R_{M}$, hence $\mathrm{dis}(m_{(h,w)},m_{k}^{\prime})\leq 2R_{M}$ and $0\leq\ell(\cdot,\cdot)\leq 2R_{M}$. Moreover, the map $\ell(\mathcal{M},\cdot)$ is $1$-Lipschitz in $\mathcal{M}^{\prime}$:
[…]
This follows from the bounded convex hull, max-hinge, and the triangle inequality.
Theorem 2 (Round Loss Reduction).
Within a fixed communication round $t$, under Assumptions 14, for any client and local iterations $e=0,\ldots,E-1$, if $\eta\in(0,2/L_{\theta})$, then
$\mathbb{E}\big[\mathcal{L}_{t,E}\big]$ $\leq\mathbb{E}\big[\mathcal{L}_{t,0}\big]-\Big(\eta-\tfrac{\eta^{2}L_{\theta}}{2}\Big)\sum_{e=0}^{E-1}\mathbb{E}\!\left[\|\nabla_{\theta}\mathcal{L}_{t,e}\|_{2}^{2}\right]$ (30)
$\quad+\frac{E\,\eta^{2}L_{\theta}}{2}\,\sigma^{2},$
and $\forall e$, $\mathbb{E}[\mathcal{L}_{t,e}]\leq 2R_{M}<\infty$.
Proof.
Assumption 4 implies $\mathrm{dis}(m,m^{\prime})\leq 2R_{M}$, hence $\max(0,\mathrm{dis}-th)\leq\max(0,2R_{M}-th)\leq 2R_{M}$, as $th>0$. Therefore $\mathbb{E}[\mathcal{L}_{t,e}]\leq 2R_{M}$.
Let the update be $\theta_{t,e+1}=\theta_{t,e}-\eta\,g_{t,e}$. By the Assumption 1, apply the one-step smooth upper bound to $\theta\mapsto\mathbb{E}[\mathcal{L}_{t,e}]$:
$\mathbb{E}[\mathcal{L}_{t,e+1}]$ $\leq\mathbb{E}[\mathcal{L}_{t,e}]+\langle\mathbb{\nabla}_{\theta}E[\mathcal{L}_{t,e}],\theta_{t,e+1}-\theta_{t,e}\rangle$
$\quad+\frac{L_{\theta}}{2}\|\theta_{t,e+1}-\theta_{t,e}\|_{2}^{2}$
@@ 5 Convergence Analysis
$\leq\mathbb{E}[\mathcal{L}_{t,e}]-\eta\,\mathbb{E}\!\left[\langle\nabla_{\theta}\mathcal{L}_{t,e},\,g_{t,e}\rangle\right]+\frac{\eta^{2}L_{\theta}}{2}\,\mathbb{E}\|g_{t,e}\|_{2}^{2}.$
By Assumption 3, $\mathbb{E}[g_{t,e}]=\mathbb{E}[\nabla_{\theta}\mathcal{L}_{t,e}]$ and $\mathbb{E}\|g_{t,e}\|_{2}^{2}=\mathbb{E}\|\nabla_{\theta}\mathcal{L}_{t,e}\|_{2}^{2}+\sigma^{2}$; substituting yields
$\mathbb{E}[\mathcal{L}_{t,e+1}]\leq\mathbb{E}[\mathcal{L}_{t,e}]-\Big(\eta-\tfrac{\eta^{2}L_{\theta}}{2}\Big)\mathbb{E}\|\nabla_{\theta}\mathcal{L}_{t,e}\|_{2}^{2}+\frac{\eta^{2}L_{\theta}}{2}\sigma^{2}.$
Summing over $e=0,\ldots,E-1$ gives (30). ∎
Theorem 3 (FedDyMem Convergence).
Under Assumptions 14 with fixed local steps $E\geq 1$, as $T\to\infty$, we have
$\frac{1}{TE}\sum_{t=1}^{T}\sum_{e=0}^{E-1}\mathbb{E}\Big[\|\nabla_{\theta}\mathcal{L}_{t,e}\|_{2}^{2}\Big]\ \longrightarrow\ 0.$ (31)
Proof.
Sum Theorem 2 over $t=1,\ldots,T$ to obtain
$\Big(\eta-\tfrac{\eta^{2}L_{\theta}}{2}\Big)\sum_{t=1}^{T}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla_{\theta}\mathcal{L}_{t,e}\|_{2}^{2}$ (32)
$\ \leq\ \sum_{t=1}^{T}\big(\mathbb{E}[\mathcal{L}_{t,0}]-\mathbb{E}[\mathcal{L}_{t,E}]\big)+T\cdot\tfrac{E\,\eta^{2}L_{\theta}}{2}\,\sigma^{2}.$
Apply a decomposition to the first term on the right:
$\sum_{t=1}^{T}\big(\mathbb{E}[\mathcal{L}_{t,0}]-\mathbb{E}[\mathcal{L}_{t,E}]\big)$ $=\mathbb{E}[\mathcal{L}_{1,0}]-\mathbb{E}[\mathcal{L}_{T,E}]$ (33)
$+\sum_{t=1}^{T-1}\Big(\mathbb{E}[\mathcal{L}_{t+1,0}]-\mathbb{E}[\mathcal{L}_{t,E}]\Big),$
Let $\Delta_{t}:=\mathbb{E}\|\mathcal{M}^{\prime\,t}-\mathcal{M}^{\prime\,t-1}\|_{2}$. By the $1$-Lipschitz property in Assumption 4,
$\mathbb{E}[\mathcal{L}_{t+1,0}]-\mathbb{E}[\mathcal{L}_{t,E}]\leq\ \Delta_{t}.$ (34)
Using the lower bound $\mathbb{E}[\mathcal{L}_{T,E}]\geq\underline{\ell}$, substitute back into (32) and divide both sides by $TE(\eta-\eta^{2}L_{\theta}/2)$ to get
@@
$\frac{1}{TE}\sum_{t=1}^{T}\sum_{e=0}^{E-1}\mathbb{E}\!\left[\|\nabla_{\theta}\mathcal{L}_{t,e}\|_{2}^{2}\right]\leq\frac{\mathbb{E}[\mathcal{L}_{1,0}]-\underline{\ell}}{(\eta-\eta^{2}L_{\theta}/2)\,TE}$ (35)
$+\frac{\eta L_{\theta}}{2(\eta-\eta^{2}L_{\theta}/2)}\,\sigma^{2}+\frac{1}{(\eta-\eta^{2}L_{\theta}/2)\,TE}\sum_{t=1}^{T-1}\Delta_{t}.$
For the first term, with $\eta\in(0,2/L_{\theta})$ and $E\geq 1$, the numerator is constant while the denominator grows linearly in $T$, hence it vanishes. For the second term, either use decreasing step sizes, leading to a noise factor proportional to $(\sum_{t,e}\eta_{t,e}^{2})/(\sum_{t,e}\eta_{t,e})\to 0$ if $\sum\eta_{t,e}=\infty$ and $\sum\eta_{t,e}^{2}<\infty$, or fix $\eta$ and increase batch size so that $\sigma^{2}\to 0$; in both cases the term vanishes. For the third term, using (16) and Assumption 4, we have
$\Delta_{t}=\alpha_{t}\,\mathbb{E}\|\widehat{\mathcal{M}}^{\prime\,t}-\mathcal{M}^{\prime\,t-1}\|\ \leq\ \frac{2R_{M}}{t+1}.$
Thus $\sum_{t=1}^{T-1}\Delta_{t}=O(\log T)$, and after normalization by $TE$ it tends to 0. Combining the three parts proves the claim. ∎
The above analysis shows that the expected loss-gradient converges to zero as iterations increase, thereby proving FedDyMem’s convergence.
6 Experiments